OpenStudy (anonymous):

please help me solve the radical function: 7-(x+6)^1/2=3 I am not sure how to make a square root on here so I used the 1/2 power

OpenStudy (anonymous):

this is to be an expression and not a function

OpenStudy (anonymous):

does this say \[7-\sqrt{x+6}=3\]?

OpenStudy (anonymous):

guess we'll never know...

OpenStudy (anonymous):

yes it does say that

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

then since \(7-4=3\) that means \[\sqrt{x+6}=4\]

OpenStudy (anonymous):

solve for \(x\) either in your head, or square and get \(x+6=16\)

OpenStudy (anonymous):

so I should come up with x=10?

OpenStudy (anonymous):

yes you should !

OpenStudy (anonymous):

okay thank you very much!

OpenStudy (anonymous):

yw